# Violation of energy conditions and the gravitational radius of the proton

### v1

@@ Abstract
Using recent lattice QCD results for gravitational form factors, we show that the Wigner-Weyl transform of the QCD energy-momentum tensor is of exotic type IV (in the Hawking-Ellis classification) in the interior of the proton, implying that in this region: (i) it has no causal eigenvectors, (ii) it violates all (linear, point-wise) energy conditions, (iii) there are local Lorentz frames in which its energy density is zero. This occurs because of the nonzero $T^{0i}$ flux associated with the angular momentum form factor $J(t)$. The energy-momentum tensor turns into a type I tensor, characteristic of “ordinary matter”, at radial distances of $1-2$ proton Compton wavelengths. This transition point defines a “gravitational radius” for the proton, whose precise determination requires accurate knowledge about QCD form factors for momentum transfer beyond $\sqrt{-t}=2$ GeV.
@@
Introduction: Gravitational form factors (GFFs) provide valuable information about the internal structure of hadrons, such as their spatial distributions of mass [1, 2], angular momentum [3], pressure, and shear forces [4, 5, 6, 7, 8, 9, 10] – for a review, see Ref. [11]. Recent measurements of such GFFs via deeply virtual Compton scattering have generated a lot of theoretical and phenomenological activity. To date, the GFFs of the proton have also been determined up to momentum transfer $\sqrt{-t}\simeq 2$ GeV from lattice QCD [12]. The GFFs parameterize the graviton-proton vertex [13, 14] and allow the construction of the energy-momentum tensor (EMT) $T_{\mu\nu}$ of the proton [4, 5] as the Wigner transform of the QCD energy-momentum operator; the Wigner transform provides a map from Hilbert space operators to quasi-probability distributions in phase space.
@@
In semiclassical gravity, this proton EMT provides a source for Einstein’s equations of general relativity (GR) for the metric of spacetime. Alternatively, instead of reading Einstein’s equations as matter determining gravity ($g_{\mu\nu}$), one can read the equation in reverse and determine how the spacetime with the desired properties determines the EMT needed to produce it. This is the point of the so-called energy conditions (ECs), used to constrain the set of possible EMTs that can source Einstein’s equations [15]. Examples are the null energy condition (NEC), the weak energy condition (WEC), the dominant energy condition (DEC), and the strong energy condition (SEC) [15]. Though not fundamental [16], ECs have been instrumental in general relativity over the years to rigorously establish, e.g., incompleteness (“singularity”) theorems, the area law theorem for black holes, and topological censorship [17, 18] Violations of ECs are widely known and, often, very consequential. For example, NEC violating EMTs can support wormholes [19]. SEC is a sufficient condition to ensure focusing of timelike or null geodesics so that gravity is always attractive in GR [15], but this condition is violated during inflation and the current accelerated expansion of the universe. Other examples where these point-wise ECs are violated, usually because of quantum effects, can be found in [20, 21, 22, 23].
@@
Note that the last expression implies a cancellation of the leading $\sim 1/t^{2}$ behavior between $A(t)$ and $2J(t)$. Besides the asymptotic scaling laws, our current theoretical knowledge of the FFs includes their determination from lattice QCD up to momentum transfer $\sqrt{-t}=2$ GeV [12], as already mentioned. Furthermore, they have also been computed in holographic models for the proton [26, 27, 28, 29]. For the electromagnetic FFs of the pion and kaon the transition to the asymptotic scaling of perturbative QCD occurs at momentum transfer of order several GeV [30].
@@
Using the parameterization (1) in (3) we may express every component of $T_{\mu\nu}(\vec{x}\,)$ in terms of the QCD form factors, see Appendix B.
@@
In terms of the EMT in (3), we may write the energy conditions in terms of the following quantities:
@@
The discriminant $\Gamma$ determines the type of energy-momentum tensor according to the Hawking-Ellis classification [15]. In regions where $\Gamma>0$, the energy-momentum tensor is of type I [15, 31], one of the eigenvectors is timelike and the others are spacelike, and all eigenvalues are real. Physically, this means that the energy flux $T^{0i}$ can always be made to vanish via a local Lorentz transformation – there is always a physical observer who finds no net energy flux in any direction – so that the EMT is diagonal in a properly defined orthonormal frame [15]. EMTs describing classical matter, where concepts such as pressure and shear stress have clear physical meaning, are expected to be of type I [15, 17]. If $\Gamma=0$, the tensor is of type II, two eigenvalues are degenerate corresponding to two null eigenvectors, and this case is typically associated with classical radiation [15]. In general, in the absence of protecting symmetries, one expects that perturbations may take the system away from type II [32]. Type III EMTs cannot occur in spherical symmetry [23], so this case is not relevant here. When $\Gamma<0$, the EMT is of type IV; one finds no causal eigenvectors, and two eigenvalues have a nonzero imaginary part. Type IV cases are exotic [20, 33, 22], and their properties away from the test-field limit in Einstein’s equations were recently investigated in [23]. For type IV EMTs, it is impossible for a physical observer not to experience a nonzero net energy flux. Furthermore, there are always local Lorentz frames in which the energy density is zero [15]. Additionally, one can show that $T_{\mu\nu}t^{\mu}t^{\nu}$ cannot be bounded from below for all unit timelike vectors $t^{\mu}$ [20]. Finally, we note that type IV EMTs violate all the point-wise ECs mentioned above, since NEC is violated.
@@
and similarly for $D(t)$ and $A(t)-2J(t)$. The parameters $\alpha_{i},\Lambda_{i}$ are given in the supplemental material of Ref. [12]. In that paper, it is shown that the dipole forms hold at least up to momentum transfer $\Delta=\sqrt{-t}=2$ GeV. However, we do know that $D(t)$ as well as $A(t)-2J(t)$ asymptotically transition into a $\sim 1/t^{3}$ fall-off. Hence, we restrict the dipole form for these two FFs to a maximum of $\Delta^{*}$; we then transition right away to the asymptotic form with a coefficient chosen so that the FFs are continuous at $\Delta^{*}$ although the first derivative will not be, of course. This is of no concern here since the energy conditions involve integrals over the FFs. We also know that $A(t)$ is supposed to fall off $\sim 1/t^{2}$ asymptotically so we assume that (11) extends to asymptotic momentum transfer in that case.
@@
At radial distances sufficiently far from the center, we find $\Gamma(r)>0$, so the tensor necessarily becomes type II before it turns into type I. For $\Delta^{*}=2$ GeV this occurs at about two proton Compton wavelengths, dropping to $1.3\bar{\lambda}$ for $\Delta^{*}=3$ GeV and to $1.1\bar{\lambda}$ for $\Delta^{*}=4$ GeV. In the far tails of the Wigner distribution, all the energy conditions mentioned above are satisfied. This can be seen from the expressions in Eq. (5): in the limit $r\to\infty$ we obtain $T_{00}\sim m$ with a positive coefficient, and $\Gamma\sim T_{00}^{2}$, $P_{t},P_{r}\to 0$.
Summary and Discussion: The form factors of the graviton-proton vertex determine the Wigner transform of the QCD energy-momentum tensor in the proton. Its spatial distribution has been used to investigate the mechanical properties of the proton, such as its pressure and shear forces. In this work, motivated by the fact that the EMT is a source term in Einstein’s equations, we have checked its properties according to the classification of Hawking-Ellis using form factors determined from lattice QCD up to momentum transfer $\sqrt{-t}=2$ GeV where we transitioned to the asymptotic forms expected from perturbative QCD. We have found that the EMT in the proton is of “ordinary” type I in the radial tails beyond a few Compton wavelengths where the standard energy conditions such as NEC, WEC, SEC, and DEC are satisfied. In the interior region, however, the proton EMT is of type IV, lacking any causal eigenvectors, and violating all the standard point-wise energy conditions. This unconventional behavior originates from the contribution of the $J$ form factor which determines the $T^{0i}$ component of the EMT. In this regard, the EMT of the pion, where $T^{0i}=0$, will not be of type IV, though the ECs may still be violated in certain regions.
@@ Appendix A Wigner transform of the QCD energy-momentum operator
$\begin{split}\int_{p}\int_{p^{\prime}}e^{-ip\cdot x}e^{ip^{\prime}\cdot x}\left<p^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p\right>&=\int\mathrm{d}^{3}P\int\frac{\mathrm{d}^{3}\Delta}{\sqrt{2p^{0}}\,(2\pi)^{3}\sqrt{2p^{\prime 0}}\,(2\pi)^{3}}\,e^{i\Delta^{0}t}e^{i(\vec{P}-\frac{1}{2}\vec{\Delta})\cdot\vec{x}}\,e^{-i(\vec{P}+\frac{1}{2}\vec{\Delta})\cdot\vec{x}}\,\left<p^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p\right>~{}.\end{split}$ (18)
@@
$\begin{split}p^{0}&=\sqrt{(\vec{P}-\vec{\Delta}/2)^{2}+m^{2}}\equiv E~{},~{}~{}~{}~{}p^{\prime 0}=\sqrt{(\vec{P}+\vec{\Delta}/2)^{2}+m^{2}}\equiv E^{\prime}~{}.\end{split}$ (20)
Note that
@@
so that
@@ Appendix B Components of the Wigner EMT in terms of the form factors
$\sqrt{p\cdot\sigma}=\sqrt{E}\left(\frac{w_{-}+w_{+}}{2}+\hat{p}\cdot\vec{\sigma}\frac{w_{-}-w_{+}}{2}\right)~{},~{}~{}\sqrt{p\cdot\bar{\sigma}}=\sqrt{E}\left(\frac{w_{-}+w_{+}}{2}+\hat{p}\cdot\vec{\sigma}\frac{w_{+}-w_{-}}{2}\right)~{},~{}~{}~{}w_{\pm}=\sqrt{1-|\vec{p}|/E}$ (36)

### v2

@@ Abstract
Using recent lattice QCD results for gravitational form factors, we show that the Wigner-Weyl transform of the QCD energy-momentum tensor is of exotic type IV (in the Hawking-Ellis classification) in the interior of the proton, implying that in this region: (i) it has no causal eigenvectors, (ii) it violates all (linear, point-wise) energy conditions, (iii) there are local Lorentz frames in which its energy density is zero. This occurs because of the nonzero $T^{0i}$ flux associated with the angular momentum form factor $J(t)$. The energy-momentum tensor turns into a type I tensor, characteristic of “ordinary matter”, at radial distances of $1-2$ proton Compton wavelengths. This transition point defines a “gravitational radius” for the proton, whose precise determination requires accurate knowledge of QCD form factors for momentum transfer beyond $-t=2$ GeV2. Imposing the validity of the averaged null energy condition (ANEC) of gravity, we derive a new model-independent inequality for the GFFs of the proton in QCD.
@@
Introduction: Gravitational form factors (GFFs) provide valuable information about the internal structure of hadrons, such as their spatial distributions of mass [1, 2], angular momentum [3], pressure, and shear forces [4, 5, 6, 7, 8, 9, 10] – for a review, see Ref. [11]. Recent measurements of such GFFs via deeply virtual Compton scattering have generated a lot of theoretical and phenomenological activity. To date, the GFFs of the proton have also been determined up to momentum transfer ${-t}=2$ GeV2 from lattice QCD [12]. The GFFs parameterize the graviton-proton vertex [13, 14] and allow the construction of the energy-momentum tensor (EMT) $T_{\mu\nu}$ of the proton [4, 5] as the Wigner transform of the QCD energy-momentum operator; the Wigner transform provides a map from Hilbert space operators to quasi-probability distributions in phase space.
@@
In semiclassical gravity, this proton EMT provides a source for Einstein’s equations of general relativity (GR) for the metric of spacetime. Alternatively, instead of reading Einstein’s equations as matter determining gravity ($g_{\mu\nu}$), one can read the equation in reverse and determine how the spacetime with the desired properties determines the EMT needed to produce it. This is the point of the so-called energy conditions (ECs), used to constrain the set of possible EMTs that can source Einstein’s equations [15]. Examples are the null energy condition (NEC), the weak energy condition (WEC), the dominant energy condition (DEC), and the strong energy condition (SEC) [15]. Though not fundamental [16], point-wise ECs have been instrumental in general relativity over the years to rigorously establish, e.g., incompleteness (“singularity”) theorems and the area law theorem for black holes [17]. Violations of ECs are widely known and, often, very consequential. SEC is a sufficient condition to ensure the focusing of timelike or null geodesics so that gravity is always attractive in GR [15], but this condition is violated during inflation and the current accelerated expansion of the universe. Other examples where point-wise ECs are violated, usually because of quantum effects, can be found in [18, 19, 20, 21]. Nevertheless, less stringent conditions obtained by imposing that ECs should only hold on average, such as the averaged null energy condition (ANEC), have played an important role in general relativity [22, 18, 23, 24, 25, 26, 27, 28] and are known to hold in unitary, Lorentz-invariant, interacting quantum field theories in Minkowski spacetime [29, 30].
@@
Note that the last expression implies a cancellation of the leading $\sim 1/t^{2}$ behavior between $A(t)$ and $2J(t)$. Besides the asymptotic scaling laws, our current theoretical knowledge of the FFs includes their determination from lattice QCD up to momentum transfer ${-t}=2$ GeV2 [12], as already mentioned. Furthermore, they have also been computed in holographic models for the proton [33, 34, 35, 36]. For the electromagnetic FFs of the pion and kaon the transition to the asymptotic scaling of perturbative QCD occurs at momentum transfer of order several GeV [37].
@@
Using the parameterization (1) in (3) we may express every component of $T_{\mu\nu}(\vec{x}\,)$ in terms of the QCD form factors, see refs. [4, 5, 9] and Appendix B.
@@
We employ the EMT of (3) to write the energy conditions in terms of the following quantities:
@@
The discriminant $\Gamma$ determines the type of energy-momentum tensor according to the Hawking-Ellis classification [15]. In regions where $\Gamma>0$, the energy-momentum tensor is of type I [15, 38], one of the eigenvectors is timelike and the others are spacelike, and all eigenvalues are real. Physically, this means that the energy flux $T^{0i}$ can always be made to vanish via a local Lorentz transformation – there is always a physical observer who finds no net energy flux in any direction – so that the EMT is diagonal in a properly defined orthonormal frame [15]. EMTs describing classical matter, where concepts such as pressure and shear stress have clear physical meaning, are expected to be of type I [15, 17]. If $\Gamma=0$, the tensor is of type II, two eigenvalues are degenerate corresponding to two null eigenvectors, and this case is typically associated with classical radiation [15]. In general, in the absence of protecting symmetries, one expects that perturbations may take the system away from type II [39]. Type III EMTs cannot occur in spherical symmetry [21], so this case is not relevant here. When $\Gamma<0$, the EMT is of type IV; one finds no causal eigenvectors, and two eigenvalues have a nonzero imaginary part. Type IV cases are exotic [18, 40, 20], and their properties away from the test-field limit in Einstein’s equations were recently investigated in [41, 21]. For type IV EMTs, it is impossible for a physical observer not to experience a nonzero net energy flux. Furthermore, there are always local Lorentz frames in which the energy density is zero [15]. Additionally, one can show that $T_{\mu\nu}t^{\mu}t^{\nu}$ cannot be bounded from below for all unit timelike vectors $t^{\mu}$ [18]. Furthermore, away from the test-field case, it has been shown [41] that there is no hypersurface-orthogonal timelike Killing vector in a spacetimes sourced by type II, III, or IV EMTs. In particular, this implies that spacetimes generated by type IV EMTs cannot be static. Finally, we note that type IV EMTs violate all the point-wise ECs mentioned above, since NEC is violated.
@@
and similarly for $D(t)$ and $A(t)-2J(t)$. The parameters $\alpha_{i},\Lambda_{i}$ are given in the supplemental material of Ref. [12]. In that paper, it is shown that the dipole forms hold at least up to momentum transfer $\Delta=\sqrt{-t}=\sqrt{2}$ GeV. However, we do know that $D(t)$ as well as $A(t)-2J(t)$ asymptotically transition into a $\sim 1/t^{3}$ fall-off. Hence, we restrict the dipole form for these two FFs to a maximum of $\Delta^{*}$; we then transition right away to the asymptotic form with a coefficient chosen so that the FFs are continuous at $\Delta^{*}$ although the first derivative will not be, of course. This is of no concern here since the energy conditions involve integrals over the FFs. We also know that $A(t)$ is supposed to fall off $\sim 1/t^{2}$ asymptotically so we assume that (11) extends to asymptotic momentum transfer in that case.
@@
At radial distances sufficiently far from the center, we find $\Gamma(r)>0$, so the tensor necessarily becomes type II before it turns into type I. For $\Delta^{*}=2$ GeV this occurs at about two proton Compton wavelengths, dropping to $1.3\bar{\lambda}$ for $\Delta^{*}=3$ GeV and to $1.1\bar{\lambda}$ for $\Delta^{*}=4$ GeV. In the far tails of the Wigner distribution, all the energy conditions mentioned above are satisfied. This can be seen from the expressions in Eq. (5): in the limit $r\to\infty$ we obtain $T_{00}\sim m/r^{3}$ with a positive coefficient, while $M^{i},P_{t},P_{r}\to 0$ faster than $r^{-3}$; hence $\Gamma\to T_{00}^{2}$.
@@
[…]
It is interesting to note that in some holographic models [35] $A=2J$, so (12) holds for $\int\mathrm{d}t\,A(t)\geq 0$ provided the mass $m\geq 0$. In QCD, the integral in (12) is finite because $A-2J\sim 1/t^{3}$ asymptotically [32], but the $-t(A-2J)$ term is not positive definite [12]. Nevertheless, we have confirmed that this inequality indeed holds for the proton EMT when the model GFFs used in our calculations of the point-wise ECs are employed. Thus, assuming that ANEC holds [^2] Besides the explicit proofs in Minkowski spacetime in [29, 30], it is known that violating ANEC would be generally catastrophic, as arguments exist stating that this could be used to build time machines [25, 26] and violate the (generalized) second law of thermodynamics [43]. Finally, we note that in Minkowski spacetime, all null lines are achronal so the ANEC condition discussed here is equivalent to the more stringent achronal averaged null energy condition [27], which is expected to hold even in curved spacetime., (12) provides a new model-independent constraint that should be satisfied by proton GFFs in QCD.
Summary and Discussion: The form factors of the graviton-proton vertex determine the Wigner transform of the QCD energy-momentum tensor in the proton. Its spatial distribution has been used to investigate the mechanical properties of the proton, such as its pressure and shear forces. In this work, motivated by the fact that the EMT is a source term in Einstein’s equations, we have checked its properties according to the classification of Hawking-Ellis using form factors determined from lattice QCD up to momentum transfer ${-t}=2$ GeV2 where we transitioned to the asymptotic forms expected from perturbative QCD. We have found that the EMT in the proton is of “ordinary” type I in the radial tails beyond a few Compton wavelengths where the standard energy conditions such as NEC, WEC, SEC, and DEC are satisfied. In the interior region, however, the proton EMT is of type IV, lacking any causal eigenvectors, and violating all the standard point-wise energy conditions. This unconventional behavior originates from the contribution of the $J$ form factor which determines the $T^{0i}$ component of the EMT. In this regard, the EMT of the pion, where $T^{0i}=0$, will not be of type IV, though the pointwise ECs may still be violated in certain regions.
The exotic properties displayed by the proton EMT, such as the fact that there are regions within the proton where its energy density can vanish or become negative, call for some caution with definitions of the proton’s mechanical properties, such as its mass radius [^3] We refer to ref. [44] for a recent comprehensive review of various proton radius definitions., using this EMT. We propose using the point where the EMT transitions from type IV to type I (thus, necessarily becoming type II at that point) to define a new “gravitational radius” for the proton, which is free from the conceptual issues mentioned above. Using current input from the lattice, we obtain that this occurs at about $1-2$ Compton wavelength. The precise value of this “classicalization” length scale above which typical mechanical properties within the proton are well-defined, standard ECs are fulfilled, and gravity is necessarily attractive, requires determining the gravitational form factors for momentum transfers beyond ${-t}=2$ GeV2, either from lattice QCD or experiments.
Finally, we showed that ANEC translates into the new model-independent constraint (12) for the proton $A(t)$ and $J(t)$ form factors in QCD. A definitive confirmation of ANEC from data reaching to sufficiently high $-t$ could motivate additional theoretical work towards its rigorous proof in asymptotically free theories. In this context, it would be interesting to also check other conditions, such as the quantum null energy condition, which has been thoroughly discussed in the literature [45, 46, 47, 48].
@@ Appendix A Wigner transform of the QCD energy-momentum operator
$\begin{split}\int_{p}\int_{p^{\prime}}e^{-ip\cdot x}e^{ip^{\prime}\cdot x}\left<p^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p\right>&=\int\mathrm{d}^{3}P\int\frac{\mathrm{d}^{3}\Delta}{\sqrt{2p^{0}}\,(2\pi)^{3}\sqrt{2p^{\prime 0}}\,(2\pi)^{3}}\,e^{i\Delta^{0}x^{0}}e^{i(\vec{P}-\frac{1}{2}\vec{\Delta})\cdot\vec{x}}\,e^{-i(\vec{P}+\frac{1}{2}\vec{\Delta})\cdot\vec{x}}\,\left<p^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p\right>~{}.\end{split}$ (19)
@@
$\begin{split}p^{0}&=\sqrt{(\vec{P}-\vec{\Delta}/2)^{2}+m^{2}}\equiv E~{},~{}~{}~{}~{}p^{\prime 0}=\sqrt{(\vec{P}+\vec{\Delta}/2)^{2}+m^{2}}\equiv E^{\prime}~{},\end{split}$ (21)
with $\Delta_{0}=E^{\prime}-E$ and $P_{0}=(E^{\prime}+E)/2$. Now we use that
@@
and interpret $x^{0},\vec{x}\left.\right|x^{0},\vec{x}\right>$ in the sense of distributions to see that its inverse may be represented as $(2\pi)^{3}\delta(\vec{P}-\vec{Q})$ when working with momentum space basis vectors. Thus, we find that
@@ Appendix B Components of the Wigner EMT in terms of the form factors
$\sqrt{p\cdot\sigma}=\sqrt{E}\left(\frac{w_{-}+w_{+}}{2}+\hat{p}\cdot\vec{\sigma}\frac{w_{-}-w_{+}}{2}\right)~{},~{}~{}\sqrt{p\cdot\bar{\sigma}}=\sqrt{E}\left(\frac{w_{-}+w_{+}}{2}+\hat{p}\cdot\vec{\sigma}\frac{w_{+}-w_{-}}{2}\right)~{},~{}~{}~{}w_{\pm}=\sqrt{1\pm|\vec{p}|/E}$ (37)
@@ Appendix C Derivation of the eigenvalues of $T_{\mu\nu}$
Appendix D Details about the ANEC calculation
The ANEC integral for the null geodesic $\ell^{\mu}=(1,n^{i})$ is
$\int\limits_{-\infty}^{\infty}\mathrm{d}\zeta\,\ell^{\mu}\ell^{\nu}\,T_{\mu\nu}=\int\limits_{-\infty}^{\infty}\mathrm{d}\zeta\,\left(T^{00}-2{n}^{i}T^{0i}+{n}^{i}{n}^{j}T^{ij}\right)~{}.$ (57)
Using the general expression in (24), this can be expressed as
$\begin{split}\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{2}}\,\frac{\delta\left(\Delta_{\mu}\ell^{\mu}\right)}{\sqrt{4EE^{\prime}}}\left[\bar{u}(p^{\prime})u(p)\,\frac{A(t)}{m}\left(\ell\cdot P\right)^{2}+\frac{J(t)}{m}\,\ell_{\mu}P^{\mu}\,\ell_{\nu}\Delta_{\rho}\,\bar{u}(p^{\prime})\,i\sigma^{\nu\rho}u(p)\right].\end{split}$ (58)
We note that the $D$ form factor does not contribute to this integral. The spinor matrix elements for general $\vec{P}$ have been computed in [49, 9],
$\begin{split}\bar{u}(p^{\prime})u(p)&={\cal N}^{-1}\left[2\left(P_{0}^{2}-\vec{P}^{2}+P_{0}m\right)+i\epsilon^{0ijk}P_{i}\Delta_{j}S_{k}\right]={\cal N}^{-1}\left[2\left(P_{0}^{2}-\vec{P}^{2}+P_{0}m\right)-i\vec{P}\cdot(\vec{\Delta}\times\vec{S})\right]~{},\\ \bar{u}(p^{\prime})i\sigma^{\mu\rho}\Delta_{\rho}u(p)&={\cal N}^{-1}\left\{P^{\mu}\Delta^{2}+m(\eta^{\mu 0}\Delta^{2}-\Delta^{\mu}\Delta^{0})+2\left[(P^{0}+m)i\epsilon^{\mu\nu\rho\lambda}\Delta_{\nu}P_{\rho}S_{\lambda}-\frac{\Delta^{2}}{4}i\epsilon^{\mu\nu\rho 0}\Delta_{\nu}S_{\rho}-(P\cdot S)i\epsilon^{\mu\nu\rho 0}\Delta_{\nu}P_{\rho}\right]\right\}~{}\end{split}$ (59)
with $S^{\mu}=(0,\vec{S})$ the rest frame spin vector; in the notation of refs. [4, 5] $\vec{S}$ had previously been denoted $\vec{\sigma}_{ss}$, up to sign. Furthermore, ${\cal N}=\sqrt{p^{0}+m}\sqrt{p^{\prime 0}+m}=\sqrt{E+m}\sqrt{E^{\prime}+m}$ and $P^{0}=(p^{0}+p^{\prime 0})/2=(E+E^{\prime})/2$. We can also write the latter for $\mu=0$ and $\mu=i$ as
$\begin{split}\bar{u}(p^{\prime})i\sigma^{0\rho}\Delta_{\rho}u(p)&={\cal N}^{-1}\left\{P^{0}\Delta^{2}-m\vec{\Delta}^{2}+2(P^{0}+m)i\vec{P}\cdot(\vec{\Delta}\times\vec{S})\right\}~{},\\ \bar{u}(p^{\prime})i\sigma^{i\rho}\Delta_{\rho}u(p)&={\cal N}^{-1}\left\{P^{i}\Delta^{2}-m\Delta^{i}\Delta^{0}+2\left[i(P^{0}+m)\left(P^{0}(\vec{\Delta}\times\vec{S})^{i}-\Delta^{0}(\vec{P}\times\vec{S})^{i}\right)+\frac{\Delta^{2}}{4}i(\vec{\Delta}\times\vec{S})^{i}-(\vec{P}\cdot\vec{S})i(\vec{\Delta}\times\vec{P})^{i}\right]\right\}~{}.\end{split}$ (60)
The ANEC integral then becomes
$\begin{split}\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{2}}\,\frac{\delta\left(\Delta_{\mu}\ell^{\mu}\right)}{\sqrt{4EE^{\prime}}}\left[\bar{u}(p^{\prime})u(p)\,\frac{A(t)}{m}\left(\ell\cdot P\right)^{2}+\frac{J(t)}{m}\,\ell_{\mu}P^{\mu}\,\ell_{\nu}\Delta_{\rho}\,\bar{u}(p^{\prime})\,i\sigma^{\nu\rho}u(p)\right].\end{split}$ (61)
The integrand can be simplified by considering the transformation $\vec{\Delta}\to-\vec{\Delta}$. This lets $\Delta^{0}\to-\Delta^{0}$ since $E\leftrightarrow E^{\prime}$, and $P^{0}\to P^{0}$. Therefore, the term in $\bar{u}(p^{\prime})u(p)$ involving $\vec{\Delta}\times\vec{S}$ is odd and drops out. Along the same lines, all terms in $\Delta_{\rho}\,\bar{u}(p^{\prime})\,i\sigma^{\nu\rho}u(p)$ which are odd under $\vec{\Delta}\to-\vec{\Delta}$ can be dropped. The ANEC integral then simplifies to
$\begin{split}\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{2}}\,\frac{\delta\left(\Delta_{\mu}\ell^{\mu}\right)}{\sqrt{4EE^{\prime}}}\left[\frac{A(t)}{m}\left(\ell\cdot P\right)^{2}\frac{2\left(P_{0}^{2}-\vec{P}^{2}+P_{0}m\right)}{\cal N}+\frac{J(t)}{m}\,\ell_{\mu}P^{\mu}\,\frac{\ell_{\nu}P^{\nu}\,\Delta^{2}+m\Delta^{2}}{\cal N}\right]~{}.\end{split}$ (62)
This is required to be $\geq 0$ for any $\vec{P}$ (and any $\vec{n}$), with $\vec{n}^{2}=1$. For $\vec{P}=0$ specifically it simplifies considerably since $E=E^{\prime}$, thus $\Delta^{0}=0,P^{0}=E,{\cal N}=E+m$:
$\begin{split}\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{2}}\,{\delta\left(\vec{n}\cdot\vec{\Delta}\right)}\left[\frac{A(t)}{m}(P^{0})^{2}+\frac{J(t)}{2m}\,\Delta^{2}\right]&=\int\frac{\mathrm{d}^{2}\Delta_{\perp}}{(2\pi)^{2}}\,\left[\left(A(-\vec{\Delta}_{\perp}^{2})-2J(-\vec{\Delta}_{\perp}^{2})\right)\frac{\vec{\Delta}_{\perp}^{2}}{4m}+mA(-\vec{\Delta}_{\perp}^{2})\right]~{}.\end{split}$ (63)
Recall that $A-2J\sim 1/\Delta_{\perp}^{6}$ asymptotically in QCD so the integral is convergent.